# Singular Value Decomposition and its Blind Spot for Quantum Chaos in Non-Hermitian Sachdev-Ye-Kitaev Models

### v2

@@ Abstract
Introduction. In recent years, the study of open quantum systems Rivas and Huelga (2012); Fazio et al. (2024) has attracted considerable attention due to the novel physical phenomena arising from interactions with the environment. Such interactions inherently render these systems non-Hermitian, complicating their theoretical description and necessitating methods beyond those developed for closed systems. Unlike standard quantum systems, non-Hermitian Hamiltonians, particularly those with broken PT symmetry, feature complex eigenvalues, posing the fundamental challenge of interpreting their physical significance Bender (2007).
@@
Acknowledgments. We would like to thank Johanna Erdmenger, Manas Kulkarni, Pratik Nandy and Yijia Zhou for valuable discussions and correspondence. MB and KBH acknowledge the support of the Foreign Young Scholars Research Fund Project (Grant No.22Z033100604). MB acknowledges the sponsorship from the Yangyang Development Fund. HSJ and JFP are supported by the Spanish MINECO ‘Centro de Excelencia Severo Ochoa’ program under grant SEV-2012-0249, the Comunidad de Madrid ‘Atracción de Talento’ program (ATCAM) grant 2020-T1/TIC-20495, the Spanish Research Agency via grants CEX2020-001007-S and PID2021-123017NB-I00, funded by MCIN/AEI/10.13039/501100011033, and ERDF ‘A way of making Europe.’ KYK was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Science, ICT &\&& Future Planning (NRF-2021R1A2C1006791) and the Al-based GIST Research Scientist Project grant funded by the GIST in 2025. KYK was also supported by the Creation of the Quantum Information Science R&\&&D Ecosystem (Grant No. 2022M3H3A106307411) through the National Research Foundation of Korea (NRF) funded by the Korean government (Ministry of Science and ICT). All authors contributed equally to this paper and should be considered as co-first authors.

### v3

@@ Abstract
Introduction. In recent years, the study of open quantum systems Rivas and Huelga (2012); Fazio et al. (2024) has attracted considerable attention due to the novel physical phenomena arising from interactions with the environment. Such interactions inherently render these systems non-Hermitian, complicating their theoretical description and necessitating methods beyond those developed for closed systems. Unlike standard quantum systems, non-Hermitian Hamiltonians, particularly those with broken parity-time (PT) symmetry, feature complex eigenvalues, posing the fundamental challenge of interpreting their physical significance Bender (2007).
@@
One might consider addressing the issue (II) above by rigidly shifting the eigenvalue spectrum to make it positive definite. In the Hermitian case, where the eigenvalues are real, this approach is indeed effective: applying the SVD method to the shifted spectrum yields the same Krylov complexity as the traditional EVD method. However, in the non-Hermitian case, this strategy fails, as the resulting Krylov complexity becomes highly sensitive to the arbitrary choice of the shift in the complex plane.
@@
Acknowledgments. We would like to thank Johanna Erdmenger, Manas Kulkarni, Pratik Nandy and Yijia Zhou for valuable discussions and correspondence. M. B. and K. B. H. acknowledge the support of the Foreign Young Scholars Research Fund Project (Grant No.22Z033100604). M. B. acknowledges the sponsorship from the Yangyang Development Fund. H. S. J. and J. F. P. are supported by the Spanish MINECO ‘Centro de Excelencia Severo Ochoa’ program under grant SEV-2012-0249, the Comunidad de Madrid ‘Atracción de Talento’ program (ATCAM) grant 2020-T1/TIC-20495, the Spanish Research Agency via grants CEX2020-001007-S and PID2021-123017NB-I00, funded by MCIN/AEI/10.13039/501100011033, and ERDF ‘A way of making Europe.’ K. Y. K. was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Science, ICT &\&& Future Planning (NRF-2021R1A2C1006791) and the Al-based GIST Research Scientist Project grant funded by the GIST in 2025. K. Y. K. was also supported by the Creation of the Quantum Information Science R&\&&D Ecosystem (Grant No. 2022M3H3A106307411) through the National Research Foundation of Korea (NRF) funded by the Korean government (Ministry of Science and ICT). K. Y. K. was also supported by the Gwangju Institute of Science and Technology (GIST) research fund (Futureleading Specialized Resarch Project, 2025) and under the framework of international cooperation program managed by the National Research Foundation of Korea (RS-2023- NR119907, FY2023).
M. B., K.-B. H., H.-S. J, X. J., K.-Y. K, and J. F. P. contributed equally to this work.